# Learning Metrics that Maximise Power for Accelerated A/B-Tests

### v1

@@ 4.4. Cost Reduction from Learnt Metrics (RQ5)
This reduction in required sample size is equal to the square of the relative $z$-score (Chapelle et al., 2012; Kharitonov et al., 2017). We visualise this quantity in Figure 5, for varying confidence levels $\alpha$, for the same Bonferroni-corrected procedure as Figure 3(b). To obtain a $z$-score for a set of metrics, we simply take the maximal score and apply a Bonferroni correction to it as laid out in Section 2.1.1. Note that this procedure depends on $\alpha$, explaining the slope in Figure 5. We observe that our learnt metrics can achieve the same level of statistical confidence as the North Star with up to 8 times fewer samples, i.e. a reduction down to $12.5\%$. This significantly reduces the cost of experimentation for the business, further strengthening the case for our learnt metrics.
Figure 5. When considering learnt metrics in conjunction with the North Star and top proxy metrics, we require significantly reduced sample sizes to obtain the same statistical confidence level as we would get from the North Star.
@@ 5. Insights from Learnt Metrics
Often, important metrics can be framed as a ratio of the means (or sums) of two existing metrics (Baweja et al., 2024; Budylin et al., 2018). Examples include click-through rate (i.e. clicks / impressions), variants of user retention (i.e. retained users / active users), or general engagement ratios (e.g. likes / video-plays). We observe that, whilst these metrics can be important from a business perspective, they typically exhibit significant type-III errors w.r.t. the North Star. Indeed, in the examples above both the numerator and denominator represent positive signals we wish to increase. Suppose an online experiment increases the number of video-plays by $Y\%$, and the overall number of likes by $X\%<Y\%$. These two positive signals will lead to a decreasing ratio, whilst we are likely to still prefer the treatment w.r.t. the North Star if $X$ and $Y$ are substantially large.
We believe this is connected to common offline ranking evaluation metrics prevalent in the recommender systems field (Steck, 2013; Jeunen, 2019). Indeed, such metrics are cumulative in nature, optimising overall value instead of some notion of value-per-item (Jeunen et al., 2023).
@@
In the previous example, we describe general count metrics for the number of likes and the number of video-play events. User behaviour on online platforms often follows a power-law distribution: a few “power users” generate the majority of such events (Chi, 2020). As a result, such metrics are easily skewed, and they are not guaranteed to accurately reflect improvements for the full population of users. Aggregating such counters per users (e.g. count the number of days a user has at least $X$ video-plays) instead of using raw event counters, provides strong and sensitive proxies to the North Star.
@@ Appendix A Additional Experimental Results
To further empirically validate our theoretical insights w.r.t. the proposed methods, we repeat the experiments reported in Section 4 on data collected for the Moj platform, and reproduce Figures 35. Results are visualised in Figure 7 Observations match our expectations, further strengthening trust in the replicability of our results.

### v2

@@ 4.4. Cost Reduction from Learnt Metrics (RQ5)
This reduction in required sample size is equal to the square of the relative $z$-score (Chapelle et al., 2012; Kharitonov et al., 2017). We visualise this quantity in Figure 5, for varying significance levels $\alpha$, for the same Bonferroni-corrected procedure as Figure 4(b). To obtain a $z$-score for a set of metrics, we simply take the maximal score and apply a Bonferroni correction to it as laid out in Section 2.1.1. Note that this procedure depends on $\alpha$, explaining the slope in Figure 5. We observe that our learnt metrics can achieve the same level of statistical confidence as the North Star with up to 8 times fewer samples, i.e. a reduction down to $12.5\%$. This significantly reduces the cost of experimentation for the business, further strengthening the case for our learnt metrics.
Figure 5. When considering learnt metrics in conjunction with the North Star and top proxy metrics, we require significantly reduced sample sizes to obtain the same statistical significance level as we would get from the North Star.
@@ 4.5. Spherical Regularisation (RQ6)
We provide source code to reproduce Figure 2 and our regularisation method at github.com/olivierjeunen/learnt-metrics-kdd-2024.
@@ 5. Insights from Learnt Metrics
Often, important metrics can be framed as a ratio of the means (or sums) of two existing metrics (Baweja et al., 2024; Budylin et al., 2018). Examples include click-through rate (i.e. clicks / impressions), variants of user retention (i.e. retained users / active users), or general engagement ratios (e.g. likes / video-plays). We observe that, whilst these metrics can be important from a business perspective, they typically exhibit significant type-III/S errors w.r.t. the North Star. Indeed, in the examples above both the numerator and denominator represent positive signals we wish to increase. Suppose an online experiment increases the number of video-plays by $Y\%$, and the overall number of likes by $X\%<Y\%$. These two positive signals will lead to a decreasing ratio, whilst we are likely to still prefer the treatment w.r.t. the North Star if $X$ and $Y$ are substantially large. Similar observations cautioning the use of ratio metrics have been made by Dmitriev et al. (2017).
We believe this is connected to common offline ranking evaluation metrics prevalent in the recommender systems field (Steck, 2013; Jeunen, 2019). Indeed, such metrics are cumulative in nature, optimising overall value instead of some notion of value-per-item (Jeunen et al., 2024).
@@
In the previous example, we describe general count metrics for the number of likes and the number of video-play events. User behaviour on online platforms often follows a power-law distribution: a few “power users” generate the majority of such events (Chi, 2020). As a result, such metrics are easily skewed, and they are not guaranteed to accurately reflect improvements for the full population of users—empirically leading to type-III/S errors w.r.t. the North Star. Aggregating such counters per users (e.g. count the number of days a user has at least $X$ video-plays) instead of using raw event counters, provides strong and sensitive proxies to the North Star.
@@ 6. Conclusions & Outlook
We believe our work opens several avenues for future work improving the efficacy of learnt metrics, by e.g. relaxing the linearity constraint we rely on. Furthermore, we wish to leverage our learnt metrics as reward signals for personalisation through machine learning models (Jeunen, 2021).
@@ Appendix A Additional Experimental Results
To further empirically validate our theoretical insights w.r.t. the proposed methods, we repeat the experiments reported in Section 4 on data collected for the Moj platform, and reproduce Figures 35. Results are visualised in Figure 7. Observations match our expectations, further strengthening trust in the replicability of our results.